OpenStudy (anonymous):

Question! (/ is a fraction bar) 5/9 divided by 3/4 = ???

OpenStudy (anonymous):

Ok I can help but first do you have any idea on what it might me?

OpenStudy (anonymous):

meaning??

OpenStudy (anonymous):

Do you think you know the answer to your problem?

OpenStudy (anonymous):

The algebraic formula for dividing fractions is a/b ÷ c/d = ad / bc. For example: 2/6 * 1/4 = 2*4 / 6*1 = 8 / 6. Reducing this fraction we get 4/3 or 1-1/3.

OpenStudy (anonymous):

So what do you think know our answer would be?

OpenStudy (anonymous):

I think it would be 20/27 what about you?

OpenStudy (anonymous):

ok let me work it out on paper first, it wont take long

OpenStudy (anonymous):

Okie Dokie!

OpenStudy (anonymous):

Came up with the same answer!

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

No problem! :D Hope I helped!! :3